OpenStudy (anonymous):

the tens digit of a certain number is 3 more than the one digit. The sum of the squares of the digits is 29. find the number.

OpenStudy (anonymous):

Let x be in the tens digit of the certain number so according to your question one's digit will be=x-3 and sum of square is 29 ie x^2 +(x-3)^2= 29 x^2+(x^2+9-6x)=29 \[2x^2 +9 -6x=29\] \[2x^2 -6x=29 -9\] \[2x^2 -6x=20\] @luhan hope you get it... can you try to solve now

OpenStudy (anonymous):

yeah, thanks :)